# Spatial-aware Transformer-GRU Framework for Enhanced Glaucoma Diagnosis from 3D OCT Imaging

### v1

@@ Abstract
Glaucoma, a leading cause of irreversible blindness, necessitates early detection for accurate and timely intervention to prevent irreversible vision loss. In this study, we present a novel deep learning framework that leverages the diagnostic value of 3D Optical Coherence Tomography (OCT) imaging for automated glaucoma detection. In this framework, we integrate a pre-trained Vision Transformer on retinal data for rich slice-wise feature extraction and a bidirectional Gated Recurrent Unit for capturing inter-slice spatial dependencies. This dual-component approach enables comprehensive analysis of local nuances and global structural integrity, crucial for accurate glaucoma diagnosis. Experimental results on a large dataset demonstrate the superior performance of the proposed method over state-of-the-art ones, achieving an F1-score of 93.58%, Matthews Correlation Coefficient (MCC) of 73.54%, and AUC of 95.24%. The framework’s ability to leverage the valuable information in 3D OCT data holds significant potential for enhancing clinical decision support systems and improving patient outcomes in glaucoma management.
@@ 1 Introduction
Figure 1: The schematic of a proposed framework for detecting glaucoma. The framework uses a large Vision Transformer (ViT-large) feature extractor called RETFound, which was pre-trained on a large OCT dataset. The symbol $\oplus$ represents the concatenation of spatial states at timestep $i$, combining information from anterior and posterior spatial locations to capture long-range spatial dependencies. The framework includes a Bidirectional Gated Recurrent Unit (GRU), Adaptive Max Pooling (AMP), and a Fully Connected (FC) layer.
@@ 2 Literature review
Numerous previous studies have focused on examining two-dimensional (2D) OCT scans, also known as B-scans, for glaucoma detection [5]. These studies have explored various deep learning methods, integrating ophthalmological domain knowledge to enhance the diagnostic efficacy of OCT imaging. For instance, Raja et al. [25] developed a deep convolutional autoencoder network to segment RNFL regions in the retina. The region-aware encoder component of their proposed model was employed for glaucoma detection, achieving an accuracy of 94.91%, F1-score of 95.77%, and an area under the curve (AUC) of 98.71% on the AFIO dataset [24]. Additionally, Garcia et al. [6] developed a hybrid neural network incorporating a VGG16 backbone tailored for processing raw OCT B-scans and an RNFL descriptor for extracting this region thickness information. Their approach achieved an accuracy of 87.88% and an F1-score of 81.82% for discriminating between healthy, early, and advanced glaucoma samples.
@@
Other imaging modalities have also been used to further improve the accuracy of glaucoma diagnosis. For instance, Song et al. [29] proposed a transformer-based architecture that uses OCT and visual field (VF) results to diagnose glaucoma. Their model incorporated deep reasoning to examine implicit pairwise relations between these two complementary modalities. The experimental results, conducted on a dataset comprising 697 glaucomatous and 698 normal scans, demonstrated an accuracy of 88.30%, an F1-score of 88.90%, and an AUC of 93.90% for glaucoma diagnosis.
@@ 3.2 Sequential Processing
To further delineate the sequential processing component of our proposed framework, the architecture employs two consecutive bidirectional GRU layers, each with a hidden size of 256. Given a sequence of feature vectors $\{f_{1},f_{2},\cdots,f_{D}\}$ where $D=64$, the bidirectional GRU processes each $f_{i}$ using Eq. 1.
@@
where $p_{t}$ is the model’s estimated probability for the class with label $t$, $\alpha_{t}$ is a weighting factor to counteract class imbalance, and $\gamma$ controls the rate at which easy examples are down-weighted. Note that the details for identifying our model’s optimal set of hyper-parameters are presented in Section 4.2.
@@ 4 Experiments
4.1 Dataset and Pre-processing
In this study, we used a dataset introduced by Maetschke et al. [20] [^1] https://doi.org/10.5281/zenodo.1481223, comprising 3D OCT scans centered on the optic nerve head. The dataset includes scans from 624 patients obtained using a Cirrus SD-OCT Scanner (Zeiss, Dublin, CA, USA), with physical dimensions of 6$\times$6$\times$2 mm and a resolution of 64$\times$64$\times$128 voxels per volume. Only scans with a signal strength of 7 or higher were included, resulting in a total of 1110 scans. These scans maintain their original laterality, avoiding any left to right eye alterations. Among these, 263 scans were classified as healthy, and 847 were diagnosed with primary open-angle glaucoma, based on glaucomatous visual field defects confirmed by at least two consecutive abnormal test results. The demographic details like gender and race distribution, along with mean values and standard deviations for patient’s age, Intraocular Pressure (IOP), Mean Field Defects (MD), and Glaucoma Hemifield Test (GHT) [2] results are presented in Table 1. It should be noted that demographic data was incomplete for some patients, and as a result, the aggregate numbers may not align exactly with the overall dataset size. For a detailed description of the dataset, you can refer to the Maestack et al [20].
@@ 4.1 Dataset and Pre-processing
In our study, we employed a 5-fold cross-validation approach to ensure our model’s robustness and generalizability across different subsets of the data. The rationale behind using 5-fold cross-validation is to provide a more reliable estimate of the model’s performance on unseen data by averaging results across multiple folds. The average number of samples allocated for training, validation, and testing in each fold is detailed in Table 2.
@@
Glaucoma 843 611 67 169
Normal 263 189 21 53
Total 1110 800 88 222
We performed a two-step pre-processing to standardize the input data for optimal integration with our proposed model. Initially, the scans underwent normalization using the mean and standard deviation of ImageNet. This step is crucial for adjusting pixel intensity values to a common scale, thereby enhancing the model’s ability to learn meaningful patterns from the data. Subsequently, the scans were resized to a uniform dimension of 64$\times$128$\times$128, ensuring consistency across the dataset and aligning with the input specifications of the pre-trained components of our model.
@@ 4.2 Implementation details and Evaluation metrics
The proposed framework is implemented using Python based on the PyTorch library version 1.8.1 on a computer powered by Intel(R) Xeon(R) Silver 4210R CPU with 16G of RAM and the NVIDIA GeForce RTX 2080 GPU with 12G of RAM. We trained the proposed framework using Adam optimizer [15], configured with the first momentum of 0.9, the second momentum of 0.999, and the stability constant of $10^{-8}$. To ensure efficient convergence, we explored the impact of varying combinations of learning rates and batch sizes on model performance across the $\{10^{-5},5\times 10^{-5},10^{-4},5\times 10^{-4},10^{-3}\}$ set for learning rates and $\{8,16,32,64,128\}$ set for batch sizes. The optimal performance on the validation set was achieved with an initial learning rate of $10^{-4}$ and a batch size of 32. We employed a learning rate scheduler to adjust the learning rate by a factor of 0.9 every 5 epochs. To address the challenge of data imbalance in our dataset, we adopted a balanced batch training strategy to ensure fair class representation and mitigate bias. The model was trained for a maximum of 100 epochs and incorporated early stopping triggered by validation loss stagnation after 6 epochs, which improved computational efficiency and prevented overfitting. We conducted an empirical exploration to find the optimal combination of five other hyper-parameters for improving the performance of the proposed glaucoma classification framework. The hyper-parameters include the hidden size of two GRU layers, the model dropout rate, and the $\alpha$ and $\gamma$ parameters in the Focal Loss function.
One important parameter for determining the capacity of the GRU architecture to learn and represent data is the hidden size of the GRU layers. Given that the architecture includes two GRU layers, we evaluated different configurations for the hidden sizes, ranging from 512, 256, to 128 on the validation set. The summary of the results is presented in Table 3. Notably, as the size of the hidden layers decreases from 512 to 128, there is a general trend toward reduced F1-scores, indicating that larger hidden layer sizes tend to capture more complex patterns and dependencies in the data. However, the optimal configuration is achieved with both GRU layers set to a hidden size of 256, resulting in the highest F1-score of 93.51%. This suggests a threshold for increasing layer size, highlighting the importance of balancing model capacity and the risk of overfitting.
@@
512 92.11 93.29 89.34
256 - 93.51 92.91
128 - - 91.24
Another parameter that helps to improve the GRU architecture generalizability is the dropout rate. As illustrated in Fig. 1(b), we investigated the effect of different dropout rates, ranging from 0.1 to 0.5, and identified 0.3 as the optimal rate. Dropout rates below this optimal threshold are associated with overfitting, where the model learns the noise in the training data too well, reducing its ability to generalize to unseen data. Conversely, dropout rates above 0.3 hinder effective learning, leading to an inadequate representation of the underlying distribution of the data.
@@
$\begin{split}\mathrm{ACC}=&~{}\frac{TP+TN}{TP+TN+FP+FN}.\\ \mathrm{SPE}=&~{}\frac{TN}{TN+FP}.\\ \mathrm{SEN}=&~{}\frac{TP}{TP+FN},\\ \mathrm{PRC}=&~{}\frac{TP}{TP+FP},\\ \mathrm{F1-score}=&~{}\frac{2\times\mathrm{SEN}\times\mathrm{PRC}}{\mathrm{SEN}+\mathrm{PRC}}.\\ \end{split}$ (5)
$\mathrm{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
where $TP$, $TN$, $FP$, and $FN$ denote true positive, true negative, false positive, and false negative, respectively. We also report area under receiver operating characteristic curve (AUC).
@@ 4.3 Experimental Results
Table 4: Summary of glaucoma detection results averaged from a 5-fold cross-validation with 95% confidence intervals (±).
Method ACC AUC SEN SPE PRC F1-score MCC
3D-CNN [20]
75.95 (± 10.43) 91.66 (± 1.98) 84.36 49.25 84.06 84.21 33.74
RETFound [35]
84.77 (± 3.66) 88.99 (± 4.02) 88.90 69.71 91.47 90.17 56.61
Proposed (ViT-large + GRU)
90.27 (± 1.44) 95.24 (± 1.69) 94.25 78.18 92.92 93.58 73.54
To evaluate the performance of our proposed framework, we first froze all the pre-trained layers within our model’s backbone feature extractors that were trained on OCT images. This strategy was employed to leverage the rich, learned representations specific to the OCT image domain, thus ensuring the extraction of precise and relevant features for glaucoma classification. We then trained our proposed framework using a 5-fold cross-validation method. The performance metrics averaged across all folds are meticulously presented in Table 4, offering a comprehensive overview of the model’s effectiveness in classifying glaucoma.
@@
As summarized in Table 4, our proposed framework significantly outperforms the other methods across all evaluation metrics. Specifically, our model achieved an accuracy of 90.27%, which is substantially higher than the 75.95% accuracy of the 3D-CNN model and the 84.77% accuracy of the standalone RETFound model. Moreover, the 3D-CNN model and standalone RETFound model exhibit wide accuracy confidence intervals of ±10.43% and ±3.66%, respectively, derived from the 5-fold cross-validation process. This indicates a lower level of reliability and variable performance across different data subsets, underscoring uncertainty in their predictive capabilities. The superior AUC of 95.24% further emphasizes the proposed model’s outstanding ability to distinguish between glaucomatous and non-glaucomatous scans.
Importantly, despite the challenges posed by an imbalanced dataset, the proposed model demonstrated a remarkable capability to accurately identify true positives and true negatives, achieving sensitivity and specificity scores of 94.25% and 78.18%, respectively. This performance is particularly noteworthy when compared to the 3D-CNN model, which showed a clear bias towards the majority class, as evidenced by its significantly lower specificity score of 49.25%. The precision, F1-score, and Matthews Correlation Coefficient (MCC) further substantiate the robustness of the proposed model, with scores that notably outperform those of the 3D-CNN and standalone RETFound models.

### v2

@@ Abstract
Glaucoma, a leading cause of irreversible blindness, necessitates early detection for accurate and timely intervention to prevent irreversible vision loss. In this study, we present a novel deep learning framework that leverages the diagnostic value of 3D Optical Coherence Tomography (OCT) imaging for automated glaucoma detection. In this framework, we integrate a pre-trained Vision Transformer on retinal data for rich slice-wise feature extraction and a bidirectional Gated Recurrent Unit for capturing inter-slice spatial dependencies. This dual-component approach enables comprehensive analysis of local nuances and global structural integrity, crucial for accurate glaucoma diagnosis. Experimental results on a large dataset demonstrate the superior performance of the proposed method over state-of-the-art ones, achieving an F1-score of 93.01%, Matthews Correlation Coefficient (MCC) of 69.33%, and AUC of 94.20%. The framework’s ability to leverage the valuable information in 3D OCT data holds significant potential for enhancing clinical decision support systems and improving patient outcomes in glaucoma management.
@@ 1 Introduction
Figure 1: Schematic of the proposed framework for detecting glaucoma. The framework uses a large Vision Transformer (ViT-large) feature extractor called RETFound, which was pre-trained on a large OCT dataset. The symbol $\oplus$ represents the concatenation of spatial states at timestep $i$, combining information from anterior and posterior spatial locations to capture long-range spatial dependencies. The framework includes a Bidirectional Gated Recurrent Unit (GRU), Adaptive Max Pooling (AMP), and a Fully Connected (FC) layer.
@@ 2 Literature review
Numerous previous studies have focused on examining two-dimensional (2D) OCT scans, also known as B-scans, for glaucoma detection [6]. These studies have explored various deep learning methods, integrating ophthalmological domain knowledge to enhance the diagnostic efficacy of OCT imaging. For instance, Raja et al. [26] developed a deep convolutional autoencoder network to segment RNFL regions in the retina. The region-aware encoder component of their proposed model was employed for glaucoma detection, achieving an accuracy of 94.91%, F1-score of 95.77%, and an area under the curve (AUC) of 98.71% on the AFIO dataset [25]. Additionally, Garcia et al. [7] developed a hybrid neural network incorporating a VGG16 backbone tailored for processing raw OCT B-scans and an RNFL descriptor for extracting thickness information in this region. Their approach achieved an accuracy of 87.88% and an F1-score of 81.82% for discriminating between healthy, early, and advanced glaucoma samples. Lee et al. [18] also used the features extracted from maps representing SD-OCT Ganglion Cell Inner Plexiform Layer (GCIPL) and RNFL thicknesses and deviations for glaucoma diagnosis. They achieved an AUC of 99%, a sensitivity of 94.70%, and a specificity of 100% when tested on a private dataset comprising 350 glaucomatous and 307 healthy SD-OCT image sets.
@@
Other data modalities have also been used to further improve the accuracy of glaucoma diagnosis. For instance, Song et al. [30] proposed a transformer-based architecture that uses OCT and visual field (VF) results to diagnose glaucoma. Their model incorporated deep reasoning to examine implicit pairwise relations between these two complementary modalities. The experimental results, conducted on a dataset comprising 697 glaucomatous and 698 normal scans, demonstrated an accuracy of 88.30%, an F1-score of 88.90%, and an AUC of 93.90% for glaucoma diagnosis.
@@ 3.3 Sequential Processing
To further delineate the sequential processing component of our proposed framework, the architecture employs two consecutive bidirectional GRU layers. Given a sequence of feature vectors $\{f_{1},f_{2},\cdots,f_{D}\}$ where $D=64$, the bidirectional GRU processes each $f_{i}$ using Eq. 1.
@@
where $p_{t}$ is the model’s estimated probability for the class with label $t$, $\alpha_{t}$ is a weighting factor to counteract class imbalance, and $\gamma$ controls the rate at which easy examples are down-weighted. Note that the details for identifying our model’s optimal set of hyper-parameters are presented in Section 4.2. The proposed glaucoma classification framework is outlined in Algorithm 1, and the implementation code is available at https://github.com/Mona-Ashtari/SpatialOCT-Glaucoma.
Input: 3D OCT volume $V\in\mathbb{R}^{D\times H\times W}$
Output: Glaucoma diagnosis Probability $p\in[0,1]$
1
2Feature Extraction:
3 Decompose $V$ into $D$ slices: $\{S_{1},S_{2},\dots,S_{D}\}$.
4 for $i=1$ to $D$ do
 $f_{i}\leftarrow\text{Pre-trained ViT-large}(S_{i})$ // Extract feature vector $f_{i}\in\mathbb{R}^{1024}$.
5
6
7Sequential Processing:
8
9Pass the sequence $\{f_{1},f_{2},\dots,f_{D}\}$ through two stacked bidirectional GRU layers:
$H_{G}\leftarrow\text{StackedGRUs}(\{f_{1},f_{2},\dots,f_{D}\})$ // Sequence of GRU outputs.
10
11Feature Pooling:
12 $h_{\mathrm{pooled}}\leftarrow\text{AdaptiveMaxPooling}(H_{G})$
13
14Classification:
$p\leftarrow\sigma(W^{\prime}\cdot h_{\mathrm{pooled}}+b^{\prime})$ // Sigmoid activation.
15
16Loss Calculation:
$\mathrm{FL}(p)=-\alpha\cdot(1-p)^{\gamma}\cdot\log(p)$ // Focal Loss.
17
18Model Training:
Optimize model parameters using the Adam optimizer with the Focal Loss.
Algorithm 1 Proposed Glaucoma Classification framework
@@ 4 Experiments
4.1 Dataset
In this study, we used a dataset introduced by Maetschke et al. [21] [^1] https://doi.org/10.5281/zenodo.1481223, comprising 3D OCT scans centered on the optic nerve head. The dataset includes scans from 624 patients obtained using a Cirrus SD-OCT Scanner (Zeiss, Dublin, CA, USA), with physical dimensions of $\times6$$\times$2 mm and a resolution of $\times64$$\times$128 voxels per volume. Only scans with a signal strength of 7 or higher were included, resulting in a total of 1110 scans. These scans maintain their original laterality, avoiding any left to right eye alterations. Among these, 263 scans were classified as healthy, and 847 were diagnosed with primary open-angle glaucoma, based on glaucomatous visual field defects confirmed by at least two consecutive abnormal test results. Figure 3 illustrates the first, middle, and last slices of 3D OCT samples from both healthy and glaucoma cases. The demographic details like gender and race distribution, along with mean values and standard deviations for patient’s age, Intraocular Pressure (IOP), Mean Field Defects (MD), and Glaucoma Hemifield Test (GHT) [3] results are presented in Table 1. It should be noted that demographic data was incomplete for some patients, and as a result, the aggregate numbers may not align exactly with the overall dataset size. For a detailed description of the dataset, you can refer to the Maestack et al [21].
@@ 4.1 Dataset
In our study, we employed a 5-fold cross-validation approach to ensure our model’s robustness and generalizability across different subsets of the data. The rationale behind using 5-fold cross-validation is to provide a more reliable estimate of the model’s performance on unseen data by averaging results across multiple folds. The average number of samples allocated for training, validation, and testing in each fold is detailed in Table 2. To ensure that the scans from the same patient are included in the same subset, the data split in each fold is performed at the subject level.
@@
Glaucoma 847 609.2 68.4 169.4
Normal 263 189.2 21.2 52.6
Total 1110 798.4 89.6 222
@@ 4.2 Implementation details and Evaluation metrics
The proposed framework is implemented using Python based on the PyTorch library version 1.8.1 on a computer powered by Intel(R) Xeon(R) Silver 4210R CPU with 16G of RAM and the NVIDIA GeForce RTX 2080 GPU with 12G of RAM. We trained the proposed framework using Adam optimizer [16], configured with the first momentum of 0.9, the second momentum of 0.999, and the stability constant of $10^{-8}$. To ensure efficient convergence, we explored the impact of varying combinations of learning rates and batch sizes on model performance across the $\{10^{-5},5\times 10^{-5},10^{-4},5\times 10^{-4},10^{-3}\}$ set for learning rates and $\{8,16,32,64,128\}$ set for batch sizes. The optimal performance on the validation set was achieved with an initial learning rate of $10^{-4}$ and a batch size of 16. We employed a learning rate scheduler to adjust the learning rate by a factor of 0.9 every 5 epochs. To address the challenge of data imbalance in our dataset, we adopted a balanced batch training strategy to ensure fair class representation and mitigate bias. The model was trained for a maximum of 100 epochs and incorporated early stopping triggered by validation loss stagnation after 6 epochs, which improved computational efficiency and prevented overfitting. We conducted an empirical exploration to find the optimal combination of five other hyper-parameters for improving the performance of the proposed glaucoma classification framework. The hyper-parameters include the hidden size of two GRU layers, the model dropout rate, and the $\alpha$ and $\gamma$ parameters in the Focal Loss function.
One important parameter for determining the capacity of the GRU architecture to learn and represent data is the hidden size of the GRU layers. Given that the architecture includes two GRU layers, we evaluated different configurations for the hidden sizes, ranging from 512, 256, to 128 on the validation set. The summary of the results is presented in Table 3. Notably, as the size of the hidden layers decreases from 512 to 128, there is a general trend toward reduced F1-scores, indicating that larger hidden layer sizes tend to capture more complex patterns and dependencies in the data. However, the optimal configuration is achieved with the first GRU layer of size 256 and the second of size 128, resulting in the highest F1-score of 93.12%. This suggests a threshold for increasing layer size, highlighting the importance of balancing model capacity and the risk of overfitting.
@@
512 92.33 92.20 91.36
256 - 91.03 93.12
128 - - 92.05
Another parameter that helps to improve the GRU architecture generalizability is the dropout rate. As illustrated in Fig. 2(b), we investigated the effect of different dropout rates, ranging from 0 to 0.5, and identified 0.3 as the optimal rate. Dropout rates below this optimal threshold are associated with overfitting, where the model learns the noise in the training data too well, reducing its ability to generalize to unseen data. Conversely, dropout rates above 0.3 hinder effective learning, leading to an inadequate representation of the underlying distribution of the data.
@@
[…]
$\mathrm{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$
where $TP$, TNTN, FPFP, and $FN$ denote true positive, true negative, false positive, and false negative, respectively. We also report area under receiver operating characteristic curve (AUC).
@@ 4.3 Experimental Results
Table 4: Summary of glaucoma detection results averaged from a 5-fold cross-validation with 95% confidence intervals (±). Values are reported as percentages, with the confusion matrix relative to the total samples per class.
Method ACC AUC SEN SPE PRC F1-score MCC Confusion Matrix
3D-CNN [21] 77.62 (± 9.78) 90.91 (± 3.34) 82.59 53.93 89.61 85.96 32.41 89.61 10.39
60.84 39.16
RETFound [36] 83.51 (± 2.26) 88.35 (± 1.73) 87.95 67.01 90.84 89.37 52.81 90.85 9.15
40.11 59.89
Proposed (ViT-large + GRU) 89.19 (± 1.89) 94.20 (± 2.00) 91.83 79.67 94.21 93.01 69.33 94.21 5.79
26.98 73.02
To evaluate the performance of our proposed framework, we first froze all the pre-trained layers within our model’s backbone feature extractors that were trained on OCT images. This strategy was employed to leverage the rich, learned representations specific to the OCT image domain, thus ensuring the extraction of precise and relevant features for glaucoma classification. We then trained our proposed framework using a 5-fold cross-validation method. The performance metrics averaged across all folds are meticulously presented in Table 4, offering a comprehensive overview of the model’s effectiveness in classifying glaucoma. Additionally, we used the attention rollout method [1] to visualize the cumulative attention across all layers of the ViT-large backbone, highlighting informative regions in OCT images (Fig. 3). The heatmaps reveal retina-specific anatomical structures, such as the optic nerve and retinal nerve fiber layer, in warmer colors, indicating their higher contribution to the extracted features.
@@
Slice 1
Slice 32
Slice 64
Slice 1
Slice 32
Slice 64
(a) Healthy
(b) Glaucoma
Figure 3: Top row: Original 3D OCT image slices. Bottom row: Corresponding attention heatmaps generated using the attention rollout method [1]. (a) Healthy scans (left 3 columns) and (b) Glaucoma scans (right 3 columns). Columns represent selected slices from the 3D OCT volume (Slice 1, Slice 32, Slice 64)
As summarized in Table 4, our proposed framework significantly outperforms the other methods across all evaluation metrics. Specifically, our model achieved an accuracy of 89.19%, which is substantially higher than the 77.62% accuracy of the 3D-CNN model and the 83.51% accuracy of the standalone RETFound model. Moreover, the 3D-CNN model and standalone RETFound model exhibit wide accuracy confidence intervals of ±9.78% and ±2.26%, respectively, derived from the 5-fold cross-validation process. This indicates a lower level of reliability and variable performance across different data subsets, underscoring uncertainty in their predictive capabilities. The superior AUC of 94.20% further emphasizes the proposed model’s outstanding ability to distinguish between glaucomatous and non-glaucomatous scans.
Importantly, despite the challenges posed by an imbalanced dataset, the proposed model demonstrated a remarkable capability to accurately identify true positives and true negatives, achieving sensitivity and specificity scores of 91.83% and 79.67%, respectively. This performance is particularly noteworthy when compared to the 3D-CNN model, which showed a clear bias towards the majority class, as evidenced by its significantly lower specificity score of 53.93%. The precision, F1-score, and Matthews Correlation Coefficient (MCC) further substantiate the robustness of the proposed model, with scores that notably outperform those of the 3D-CNN and standalone RETFound models.